f(x) = 3(x-1)(x+2)/ (x-4) (x+2)

a) state vertical asymptotes

b) state the roots


1
Expert's answer
2021-09-14T06:05:38-0400

ANSWER. a)The line "x=4" is the vertical asymptote

b) "x=1" is the only root of the function.

EXPLANATION

The function "f" is a rational function since it has the form "f(x)=\frac { P(x) }{ Q(x) }" where "P" and "Q" are polynomials. "P(x)=3(x-1)(x+2), Q(x)=(x-4)(x+2)" .The domain of definition of function "f" is the set "D=\left( -\infty ,+\infty \right) \diagdown \left\{ x:Q(x)=0 \right\} =\left( -\infty ,-2 \right) \cup \left( -2,4 \right) \cup \left( 4,+\infty \right) \ ." If "x\in D" then "f(x)=\frac { 3(x-1)\ }{ (x-4)\ } ." In the domain of the definition of the function "f" is continuous. Since "\lim _{ x\rightarrow -2 }{ f(x)=\lim _{ x\rightarrow -2 }{ \frac { 3(x-1)\quad }{ (x-4)\quad } = } \frac { 3\cdot (-2-1) }{ (-2-4) } =\frac { 3 }{ 2 } }" , then "x=-2" point of removable of discontinuity . a)The line "x=4" is the vertical asymptote, because "\lim _{ x\rightarrow 4 }{ f(x)=\lim _{ x\rightarrow 4 }{ \frac { 3(x-1)\quad }{ (x-4)\quad } =\infty } \ }."

b) "f(x)=0" if and only if "x=1." So, "x=1" is the only root of the function.